OpenStudy (anonymous):

Help Solve. 2(a + 4) + 6a = 48 a = _________ Question 5.5. The width of a rectangle is 8 cm less than its length. The perimeter of the rectangle is 52 cm. What is the width of the rectangle? (Points : 1) 7 cm 9 cm 15 cm 17 cm

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

2(a + 4) + 6a = 48 Distribute: 2a + 8 + 6a = 48 Subtract 8 to both sides: 2a + 6a = 40 Add: 8a = 40 Divide 8 to both sides: a = 5

OpenStudy (igreen):

@Xgloves11

OpenStudy (anonymous):

none of the answers are five

OpenStudy (anonymous):

if you look in your book it will tell you something

OpenStudy (igreen):

What are your answer choices for your 1st question? @Xgloves11

OpenStudy (anonymous):

A.7 B.9 C.15 D.17

OpenStudy (igreen):

For the 2nd question, plug in what you know from the Perimeter formula. P = 2L + 2W Where P = Perimeter, L = Length, and W = Width. 52 = 2(8) + 2W 52 = 16 + 2W 36 = 2W w = 18

OpenStudy (anonymous):

oh thx

OpenStudy (anonymous):

2(a + 4) + 6a = 48 Distribute. 2a + 2(4) + 6a = 48

OpenStudy (anonymous):

Simplify now. 2a + 8 + 6a = 48 Now combine like terms. Do you know how to do that?

OpenStudy (anonymous):

yes i think it is A or B

OpenStudy (anonymous):

6a + 2a = 8a So: 8a + 8 = 48. After this, I think you subtract eight and then divide to get your answer(: try it and see what happens.

OpenStudy (anonymous):

@iGreen wow I got the same answer as you...

OpenStudy (anonymous):

me too its 5

OpenStudy (anonymous):

okay great! So I got the right answer(:

OpenStudy (anonymous):

Did my answer help you in any way?

OpenStudy (anonymous):

yea it was right thx

OpenStudy (anonymous):

no problem(: I have a question, what are medals for? I'm new so this is a stupid question but ... like everyone wants them or something?

OpenStudy (igreen):

You give medals to people who have helped you answer your question.

OpenStudy (anonymous):

gotcha

OpenStudy (anonymous):

its basically points to get an award like get 500 medals and you can become a mathlete @AubzQuintanilla

OpenStudy (igreen):

Lol like me :P You also need 500 smartcents, which you get from medals.

OpenStudy (anonymous):

wow